OpenStudy (anonymous):

Use the quadratic formula to solve the equation. -4x^2-3x+2=0

OpenStudy (michele_laino):

we have to rewrite your equation like this: \[4{x^2} + 3x - 2 = 0\]

OpenStudy (michele_laino):

after that, we have to apply this formula: \[x = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] where \(a=4,b=3,c=-2\) Please try

OpenStudy (anonymous):

\[x=\frac{ -3\pm \sqrt{-23} }{8}\] So would it look like this?

OpenStudy (michele_laino):

I got this: \[x = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}} = \frac{{ - 3 \pm \sqrt {9 + 32} }}{8} = \frac{{ - 3 \pm \sqrt {41} }}{8}\]

OpenStudy (anonymous):

So it looks like I messed up some how in that one section, so can you explain how you got that?

OpenStudy (michele_laino):

step by step: \[\sqrt {{b^2} - 4ac} = \sqrt {{3^2} - 4 \cdot 4 \cdot \left( { - 2} \right)} = \sqrt {9 + 32} = \sqrt {41} \]

OpenStudy (anonymous):

oh I see what I did wrong okay minor mistake. Okay so what next do I do?

OpenStudy (michele_laino):

so, we have these solutions: \[{x_1} = \frac{{ - 3 + \sqrt {41} }}{8},\quad {x_2} = \frac{{ - 3 - \sqrt {41} }}{8}\]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

:)